A 8.2 resistor and a 5.9 resistor are connected in parallel with a battery. The current across the 8.2 resistor is measured as 0.25 A. Find the potential difference across the battery. V
Answer #1

here,

R1 = 8.2 ohm

R2 = 5.9 ohm

the current across 8.2 ohm , I1 = 0.25 A

as the resistances are connected in parallel with the battery

the potential difference across the battery and resistors will be same

the potential difference across the batery , V = potential difference across 8.2 ohm

V = I1 * R1

V = 8.2 * 0.25 V = 2.05 V